Rachel drives at a rate of 40 miles per hour for 3 hours on a country road and then drives for 2 hours at a rate of 55 miles per hour on a highway. What was her average speed, in miles per hour, for the entire trip?
(A)45 (B) 46 (C) 47.5 (D) 50 (E) 51

Answer by mananth(16949)

40 mph for 3 hours.
she travels 120 miles
.
.
55mph for 2 hours
she travels 110 miles
.
Total distance traveled = 120+110 = 230 miles
total time taken = 5 hours
Average speed for total journey = 230/5
= 46 miles/hr